# Neural ODEs for holographic transport models without translation symmetry

### v2

@@ Abstract
We investigate the data-driven holographic transport models without translation symmetry. Our data are the real part of the frequency-dependent shear viscosity, denoted as $\eta_{\mathrm{re}}(\omega)$. We develop a radial flow equation of the shear response and establish its relation to $\eta_{\mathrm{re}}(\omega)$ for a wide class of holographic models. This allows us to determine $\eta_{\mathrm{re}}(\omega)$ of a strongly coupled field theory by the black hole metric and the graviton mass. The latter serves as the bulk dual to the translation symmetry breaking on the boundary. We convert the flow equation to a Neural Ordinary Differential Equation (Neural ODE), which is a neural network with continuous depth and produces output through a black-box ODE solver. We test the Neural ODE in three well-known holographic models without translation symmetry. Given either the metric or the mass, we display that the Neural ODE can learn the other with high accuracy. We also illustrate that the learned metric can be used to predict the derivative of entanglement entropy $S$ with respect to the size of entangling region $l$, which suggests an intriguing relationship between $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$.
@@ 1 Introduction
To address these discretization problems, Neural ODEs have been introduced for AdS/DL Hashimoto:2020 ; Hashimoto:2022eij . This is a relatively new family of deep neural network models that generalize the standard layer-to-layer propagation to continuous-depth models Chen:2018 . In essence, Neural ODEs employ a neural network to parameterize the derivative of the hidden state, while the output is computed using a black-box adaptive differential equation solver. The application of Neural ODEs in holographic modeling, as highlighted in Hashimoto:2020 , showcases the potential of AI as more than just a data processing tool in theoretical physics, but also as a valuable aid in scientific discoveries.
@@
The purpose of this paper is twofold. On a technical level, we aim to replace the discrete architecture in Wu2020 with Neural ODEs. Note that the metric ansatz was limited to a polynomial in Hashimoto:2020 ; Hashimoto:2022eij , which constrains the representation capability of Neural ODEs. In this work, we will eliminate this constraint by incorporating the neural network into the ansatz. It should be pointed out that switching from polynomials to neural networks increases the training difficulty significantly. Fortunately, we find an effective training method. The key strategies include using three optimizers (RMSProp, Adam, BFGS) in sequence, as well as employing a larger number of initial values and the early-stopping technique to prevent Neural ODEs from getting stuck with bad parameters. On a physical level, we intend to break the translation symmetry that is respected by the theoretical framework in Wu2020 . The latter is indispensable to describe the real-world condensed matter. However, the breaking of translation symmetry presents an important challenge. Following the spirit of AdS/DL, one would expect the framework of the theory to be as universal as possible, so that the machine has the potential to learn novel metrics from the data of field theories, where the gravity dual may not be explicit. In Hashimoto:2018ftp , the universality is achieved under the probe limit, where the bulk action generating the background metric can be entirely unknown. In Wu2020 , the probe limit is not invoked, but the universality depends heavily on the relationship between the shear viscosity of the boundary field theory and the shear response on the UV boundary. If the translation symmetry is broken, this relationship usually no longer holds. In this work, through a careful analysis of the UV behavior related to the holographic renormalization Skenderis0209 ; Papadimitriou2016 , we find that the shear viscosity and the shear response on the boundary usually have the same real part, even if the translation symmetry is broken. This crucial observation motivates us to use only the real part of the shear viscosity as the training data.
@@ 2 Holographic flow equation
Importantly, eq. (1) and eq. (2) are homogeneous and isotropic along the field theory directions, although we have not assumed that the matter fields are homogeneous. When the black hole is perturbed by the time-dependent shear mode $\delta g_{xy}=g_{xx}h(r)e^{-i\omega t}$, the wave equation has a general form Hartnoll1601  [^2] Here we have assumed that this mode decouples from other perturbations. Since $\delta g_{xy}$ transforms as a parity-even tensor mode in $x-y$ plane, this assumption is quite general. However, one can find a counterexample in the Einstein-$SU(2)$ theory Son1311 , where the coupling appears from a particular background that locks the symmetries in spacetime and $SU(2)$ vector space.
@@ 4.3 Generalized axions
$f(z)$ $=$ $1-z^{3}+\frac{1}{2}\sum_{n=1,5}\gamma_{n}\frac{z^{3}-z^{2n}}{3-2n},$ (41)
@@
Using eq. (41) and eq. (45), the data are generated in the left panel of figure 4. Here we have fixed the parameters $\gamma_{1}=\gamma_{5}=1$. We exhibit the training results in middle and right panels of figure 4, as well as in figure 6 and table 1.
@@ 5 Conclusion and discussion
Although some assumptions still remain, the current theoretical framework, which mainly consists of the wave equation (3), the regularity at the horizon (7), and the boundary relation (8), is very general. Considering that the holographic calculation of shear viscosity does not impose the probe limit, this generality is quite rare.
@@
In table 1, it is shown that the Neural ODE can achieve high accuracy for the bulk reconstruction. However, we caution that this is associated with the current model parameters ($\alpha,\beta,\gamma$), which are simply fixed to 1. As these parameters increase, the temperature of black holes decreases. Meanwhile, we observe a rapid degradation in the performance of Neural ODEs [^5] Similar phenomena were previously reported in Li:2022zjc .. To address this issue, we will input the temperature and several inductive biases, the most important of which is to change the loss function. In this way, Neural ODEs can still perform well, see Appendix C.
@@
In Hashimoto:2018bnb , after the bulk metric is learned from lattice QCD data, the quark-antiquark potential is calculated using the holographic method. Intriguingly, this potential exhibits both a linear confining part and a Debye screening part. In Appendix D, we conduct a similar exploration.
Using the bulk metric $f(z)$ learned from the shear viscosity $\eta_{\mathrm{re}}(\omega)$, we calculate the derivative $S^{\prime}(l)$, where $S$ is the holographic EE and $l$ is the spatial size of entangling region [^6] We focus on the derivative of EE rather than the EE itself because the former is not sensitive to the UV cutoff Jokela:2020auu ; Jokela2304 ; XWB2023 .. Across a wide range of $l$ spanning several orders of magnitude, we find that the MRE of $S^{\prime}(l)$ is very small. This indicates that there is an intriguing relationship between $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$. Note that this relationship may be very complex from the perspective of field theories, partly because $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$ belong to two distinct classes of observables: quantum information measures and transport coefficients. Meanwhile, $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$ are mediated by the bulk metric, connected via the Neural ODE and the Ryu-Takayanagi (RT) formula Ryu:2006bv , see figure 5. More importantly, the relationship between $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$ is independent with the probe limit and is not restricted to a specific model. Therefore, we suspect that it may appear in some real-world strongly coupled systems. In particular, given the recent progress in lattice gauge theories with respect to both $S^{\prime}(l)$ and the spectral function related to shear viscosity Jokela2304 ; Altenkort2211 ; Yao2402 , we expect that this conjecture can be verified in the near future.
Figure 5: The schematic diagram for the relationship between $\eta_{\mathrm{re}}(\omega)$ and $S^{\prime}(l)$. At the gravity level, they are connected by the Neural ODE and the RT formula through the bulk metric $f(z)$, which is represented by the straight arrows above. The winding arrow below implies that their relationship may be very complex from the perspective of field theories.
@@
The current theoretical framework necessitates that the bulk spacetime be asymptotically AdS. However, we deliberately refrain from explicitly communicating this information to the machine. This intentional omission increases the training difficulty considerably, but the performance of Neural ODEs remains excellent. This suggests the potential for developing a machine-learning framework that can extend beyond asymptotic AdS, capable of accommodating the Lifshitz and hyperscaling-violated geometries, as well as other novel behaviors [^7] By constructing a free bulk theory from the generalized free fields on the boundary, it was found recently that the Sachdev-Ye-Kitaev model retains geometric bulk features beyond its conformal limit Qi2306 . Notably, the curvature of the emergent spacetime diverges near the boundary.. Ultimately, this work falls within the realm of reverse engineering of AdS/CFT, where the holographic dictionary is given a priori. One can also develop holographic machine learning that is independent of AdS/CFT You:2017guh ; Hu:2019nea ; Han:2019wue ; Lam:2021ugb , which explores the emergence of spacetime and gravity from a more fundamental level.
Note added: Following the completion of this work, we became aware of ref. Ahn2401 , where Neural ODEs are used to learn from optical conductivity based on the linear axion model. Encouragingly, it is illustrated that the bulk metric can be reconstructed from the data of the real material UPd2Al3.
@@ Appendix A From shear response to shear viscosity
Interestingly, it is independent with the energy-momentum tensor and the graviton mass. In contrast, we cannot write down the general form of $L_{\mathrm{matter}}^{(2)}$.
@@
The counterterms of the linear axion model are simple Andrade:1311 :
@@
In terms of the holographic renormalization studied in Ma2208 , one can infer that the counterterms of the current generalized axion model are given by
@@
From (A.14), (A.16), and (A.18), it is obvious that they can only generate real $L_{\mathrm{matter}}^{(3)}$.
@@
where $\left(\Psi,\partial\Psi\right)_{t\cdots}$ denotes certain matter field or its spatial derivative with the indices $t\cdots$. The expression of $L_{\mathrm{im}}$ is understood that the subscript $t$ in $\partial_{t}\gamma_{xy}$ must be paired with the subscript $t$ in $\left(\Psi,\partial\Psi\right)_{t\cdots}$ in order to contract with the superscripts of $g^{tt}$. Obviously, $\Psi$ is neither a scalar nor a spinor since they have no spacetime indices. $\Psi$ is also not a U(1) vector: Due to the gauge symmetry, the U(1) vector can only appear in the form of $F_{ij}F^{ij}$, which cannot be paired with $\partial_{t}\gamma_{xy}$. In light of these analysis, one can find that our assumption is valid for all Einstein-Maxwell-Dirac-scalar models with parity symmetry and neutral matters. For more general models, we cannot provide a similar proof by checking eq. (A.19). Nevertheless, keeping in mind the intrinsic structure of counterterms [^9] Even if one can construct a counterterm ansatz that yields $L_{\mathrm{im}}$, this does not necessarily imply that it is a genuine counterterm for a specific model. In Ma2208 , it was found that the exact ansatz of counterterms can be generated from the radial Hamiltonian, indicating the existence of an intrinsic structure of counterterms., it is reasonable to suspect that the current assumption may be valid in a more general situation [^10] For example, consider the famous model of holographic superconductor Hartnoll0803 . The charged complex scalar field $\Phi$ is coupled with the U(1) vector $A$, which in principle might produce: $L_{\mathrm{im}}\sim\gamma^{tt}\partial_{t}\gamma_{xy}\gamma_{xy}\left|(\partial_{t}-iA_{t})\Phi\right|$. However, the matter contribution to the counterterm of this model is just $L_{\mathrm{matter}}^{(1)}\sim\left|\Phi\right|^{2}$.. In fact, we are not aware of any obvious counterexamples.
To proceed, we extract the retarded correlation from $S_{\mathrm{ren}}$:
$G_{T^{xy}T^{xy}}^{R}(\omega)=L^{(3)}+3\frac{h^{(3)}}{h^{(0)}},$ (A.20)
@@ Appendix B Training scheme and report
There are various hyperparameter tuning methods, which consume different computing resources. Here we propose a simple way to quickly find the exponent $b$ in the square of graviton mass. We define the search range as $[0,4]$ with the step size 1. We use the standard cross-validation method with the simple hold-out technique Arlot . We separate the dataset with the frequency range $[0.01,\omega_{\max}]$ into two parts. The part with $[0.01,\omega_{\max}-1]$ goes to the training set and the part with $[\omega_{\max}-0.99,\omega_{\max}]$ to the validation set. The rationale for this separation lies in the dependence of the exponent $b$ on the UV physics. We use the optimizers RMSProp and Adam in order. Their learning rates, epoch numbers, and batch sizes are set as (0.001, 0.0001), (10, 10) and (1, 1). We also add the L1 regularization term $L_{1}=0.1\times\left|\theta\right|$ in the loss function. For each $b$ in the search range, we train 5 times and take the average validation error to select the optimal $b$. As long as $\omega_{\max}$ is large enough, the correct $b$ can be found for the current three targets. Specifically, $\omega_{\max}=2$ is large enough for the MG and LA models, and $\omega_{\max}=4$ is large enough for the GA model, see figure 6.
@@
Figure 6: The root mean square error (RMSE) for evaluating the hyperparameter $b$ in the MG (left), LA (middle) and GA (right) models. Each $b$ is trained five times, yielding five nearly coincident RMSE points in most cases. As one can see, the optimal hyperparameters for three models are 1, 2, and 2, respectively.

### v3

@@ Abstract
We investigate the data-driven holographic transport models without translation symmetry, focusing on the real part of frequency-dependent shear viscosity, $\eta_{\mathrm{re}}(\omega)$. We develop a radial flow equation of the shear response and establish its relation to $\eta_{\mathrm{re}}(\omega)$ for a wide class of holographic models. This allows us to determine $\eta_{\mathrm{re}}(\omega)$ of a strongly coupled field theory by the black hole metric and the graviton mass. The latter serves as the bulk dual to the translation symmetry breaking on the boundary. We convert the flow equation to a Neural Ordinary Differential Equation (Neural ODE), which is a neural network with continuous depth and produces output through a black-box ODE solver. Testing the Neural ODE on three well-known holographic models without translation symmetry, we demonstrate its ability to accurately learn either the metric or mass when given the other. Additionally, we illustrate that the learned metric can be used to predict the derivative of entanglement entropy $S$ with respect to the size of entangling region $l$.
@@ 1 Introduction
To address these discretization problems, Neural ODEs have been introduced for AdS/DL Hashimoto:2020 ; Hashimoto:2022eij ; Ahn2401 . This is a relatively new family of deep neural network models that generalize the standard layer-to-layer propagation to continuous-depth models Chen:2018 . In essence, Neural ODEs employ a neural network to parameterize the derivative of the hidden state, while the output is computed using a black-box adaptive differential equation solver. The application of Neural ODEs in holographic modeling, as highlighted in Hashimoto:2020 , showcases the potential of AI as more than just a data processing tool in theoretical physics, but also as a valuable aid in scientific discoveries.
@@
The purpose of this paper is twofold. On a technical level, we aim to replace the discrete architecture in Wu2020 with Neural ODEs. Note that the metric ansatz was limited to a polynomial in Hashimoto:2020 ; Hashimoto:2022eij ; Ahn2401 , which constrains the representation capability of Neural ODEs. In this work, we will eliminate this constraint by incorporating the neural network into the ansatz. It should be pointed out that switching from polynomials to neural networks increases the training difficulty significantly. Fortunately, we find an effective training method. The key strategies include using multiple optimizers in sequence, as well as employing a larger number of initial values and the early-stopping technique to prevent Neural ODEs from getting stuck with bad parameters. On a physical level, we intend to break the translation symmetry that is respected by the theoretical framework in Wu2020 . The latter is indispensable to describe the real-world condensed matter. However, the breaking of translation symmetry presents an important challenge. Following the spirit of AdS/DL, one would expect the framework of the theory to be as universal as possible, so that the machine has the potential to learn novel metrics from the data of field theories, where the gravity dual may not be explicit. In Hashimoto:2018ftp , the universality is achieved under the probe limit, where the bulk action generating the background metric can be entirely unknown. In Wu2020 , the probe limit is not invoked, but the universality depends heavily on the relationship between the shear viscosity of the boundary field theory and the shear response on the UV boundary. If the translation symmetry is broken, this relationship usually no longer holds. In this work, through a careful analysis of the UV behavior related to the holographic renormalization Skenderis0209 ; Papadimitriou2016 , we find that the shear viscosity and the shear response on the boundary usually have the same real part, even if the translation symmetry is broken. This crucial observation motivates us to use only the real part of the shear viscosity as the training data [^1] It is instructive to note that the shear viscosity does not have the hydrodynamic interpretation when the translation symmetry is broken, but it still quantifies the entropy production rate due to the strain Hartnoll1601 . In fact, the definition of the viscosity tensor is independent with the hydrodynamics: it is the linear response of the stress tensor under a time–dependent deformation Read1207 . Recent studies on the viscosity for the systems without translation symmetry can be found, e.g., in Burmistrov1901 ; Baggioli1910 ..
@@ 2 Holographic flow equation
Importantly, eq. (1) and eq. (2) are homogeneous and isotropic along the field theory directions, although we have not assumed that the matter fields are homogeneous. When the black hole is perturbed by the time-dependent shear mode $\delta g_{xy}=g_{xx}h(r)e^{-i\omega t}$, the wave equation has a general form Hartnoll1601  [^3] Here we have assumed that this mode decouples from other perturbations. Since $\delta g_{xy}$ transforms as a parity-even tensor mode under the rotation in $x-y$ plane, this assumption is quite general. However, one can find a counterexample in the Einstein-$SU(2)$ theory Son1311 , where the coupling appears from a particular background that locks the symmetries in spacetime and $SU(2)$ vector space.
@@ 4.3 Generalized axions
$f(z)$ $=$ $1-z^{3}+\frac{1}{2}\sum_{n=1,5}\gamma_{n}\frac{z^{3}-z^{2n}}{3-2n},$
@@
Using eq. (4.3) and eq. (44), the data are generated in the left panel of figure 4. Here we have fixed the parameters $\gamma_{1}=\gamma_{5}=1$. We exhibit the training results in middle and right panels of figure 4, as well as in figure 6 and table 1.
@@ 5 Conclusion and discussion
Although some assumptions still remain, the current theoretical framework, which mainly consists of the wave equation (3), the regularity at the horizon (7), and the boundary relation (8), is very general. Considering that the holographic calculation of shear viscosity does not impose the probe limit, this generality is quite rare and precious.
@@
In Table 1, it is shown that the Neural ODE can achieve high accuracy for the bulk reconstruction. However, we caution that this is associated with the current model parameters ($\alpha,\beta,\gamma$), which are simply fixed to 1. As these parameters increase, the temperature of black holes decreases. Meanwhile, we observe a rapid degradation in the performance of Neural ODEs [^6] Similar phenomena were previously reported in Li:2022zjc .. To address this issue, we will specify the temperature by hyperparameter tuning and input several inductive biases. In this way, Neural ODEs can still perform well, see Appendix C.
@@
In Hashimoto:2018bnb , after the bulk metric is learned from lattice QCD data, the quark-antiquark potential is calculated using the holographic method. Interestingly, this potential exhibits both a linear confining part and a Debye screening part. In Appendix D, we conduct a similar exploration.
Using the bulk metric $f(z)$ learned from the shear viscosity $\eta_{\mathrm{re}}(\omega)$, we calculate the derivative $S^{\prime}(l)$, where $S$ is the holographic EE and $l$ is the spatial size of entangling region [^7] We focus on the derivative of EE rather than the EE itself because the former is not sensitive to the UV cutoff Jokela:2020auu ; Jokela2304 ; XWB2023 .. Across a wide range of $l$ spanning several orders of magnitude, we find that the MRE of $S^{\prime}(l)$ is very small. This indicates that we have established a relationship between $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$. Note that this relationship may be very complex from the perspective of field theories, partly because $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$ belong to two distinct classes of observables: quantum information measures and transport coefficients. Nevertheless, using the bulk metric as a medium, $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$ can indeed be linked via the Neural ODE and the Ryu-Takayanagi (RT) formula Ryu:2006bv , see figure 5.
Figure 5: The schematic diagram for the relationship between $\eta_{\mathrm{re}}(\omega)$ and $S^{\prime}(l)$. At the bulk level, they are connected by the Neural ODE and the RT formula through the metric, which is represented by the straight arrows above. The curved arrow below indicates that the relationship between them can be very complicated from the perspective of field theory.
More importantly, the relationship between $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$ is independent with the probe limit and is not confined to a specific model [^8] For any effective methods of bulk metric reconstruction, their field-theory observables are interrelated, reflecting at least a redundancy in the bulk geometry information. However, each reconstruction method has its applicable range. If these ranges are very limited or only slightly overlap, we would not expect a clear relationship between the field-theory observables.. Based on its generality and robustness [^9] It is found that the relative error of $S^{\prime}(l)$ decreases oscillatingly as $l$ increases, indicating a more robust relationship in the IR. This robustness could facilitate the emergence of the relationship in the real-world strongly coupled systems that lack conformal symmetry in the UV., we conjecture that this relationship may manifest in some real-world strongly coupled systems. In particular, given the recent progress in lattice gauge theories with respect to both $S^{\prime}(l)$ and the spectral function related to shear viscosity Jokela2304 ; Altenkort2211 ; Yao2402 , we expect that this conjecture could be verified in the near future.
@@
The current theoretical framework necessitates that the bulk spacetime be asymptotically AdS. However, we deliberately refrain from explicitly communicating this information to the machine. Our intentional omission increases the training difficulty considerably, but the performance of Neural ODEs remains excellent. This encouraging result suggests the potential to develop a machine-learning framework that extends beyond asymptotic AdS, encompassing the Lifshitz and hyperscaling-violated geometries, as well as other novel behaviors [^10] By constructing a free bulk theory from the generalized free fields on the boundary, it was found recently that the Sachdev-Ye-Kitaev model retains geometric bulk features beyond its conformal limit Qi2306 . Notably, the curvature of the emergent spacetime diverges near the boundary..
Ultimately, this work falls within the realm of reverse engineering of AdS/CFT, where the holographic dictionary is given a priori. One can also develop holographic machine learning that is independent of AdS/CFT You:2017guh ; Hu:2019nea ; Han:2019wue ; Lam:2021ugb , which explores the emergence of spacetime and gravity from a more fundamental level.
@@ Appendix A From shear response to shear viscosity
Interestingly, it is independent with the energy-momentum tensor and the gravition mass. In contrast, we can not write down the general form of $L_{\mathrm{matter}}^{(2)}$.
@@
The counterterms of the linear axion model are simple:
@@
In terms of the holographic renormalization studied in Ma2208 , one can infer that the counterterms of generalized axion model are given by
@@
From eq. (A.14), eq. (A.16), and eq. (A.18), it is obvious that they can only generate real $L_{\mathrm{matter}}^{(3)}$.
@@
where $\left(\Psi,\partial\Psi\right)_{t\cdots}$ denotes certain matter field or its spatial derivative with the indices $t\cdots$. The expression of $L_{\mathrm{im}}$ is understood that the subscript $t$ in $\partial_{t}\gamma_{xy}$ must be paired with the subscript $t$ in $\left(\Psi,\partial\Psi\right)_{t\cdots}$ in order to contract with the superscripts of $g^{tt}$. Obviously, $\Psi$ is neither a scalar nor a spinor since they have no spacetime indices. $\Psi$ is also not a U(1) vector: Due to the gauge symmetry, the U(1) vector can only appear in the form of $F_{ij}F^{ij}$, which cannot be paired with $\partial_{t}\gamma_{xy}$. In light of these analysis, one can find that our assumption is valid for all Einstein-Maxwell-Dirac-scalar models with parity symmetry and neutral matters. For more general models, we cannot provide a similar proof by checking eq. (A.19). Nevertheless, keeping in mind the intrinsic structure of counterterms [^12] Even if one can construct a counterterm ansatz that yields $L_{\mathrm{im}}$, this does not necessarily imply that it is a genuine counterterm for a specific model. In Ma2208 , it was found that the exact ansatz of counterterms can be generated from the radial Hamiltonian, indicating the existence of an intrinsic structure of counterterms., it is reasonable to suspect that the current assumption may be valid in a more general situation [^13] For example, consider the famous model of holographic superconductor Bradlyn0803. The charged complex scalar field $\Phi$ is coupled with the U(1) vector $A$, which in principle might produce: $L_{\mathrm{im}}\sim\gamma^{tt}\partial_{t}\gamma_{xy}\gamma_{xy}\left|(\partial_{t}-iA_{t})\Phi\right|$. However, the matter contribution to the counterterm of this model is just $L_{\mathrm{matter}}^{(1)}\sim\left|\Phi\right|^{2}$.. In fact, we are not aware of any obvious counterexamples.
To proceed, we consider a field theory lived in a two-dimensional flat space and define the retarded correlatorWu2020 ; Wu2015
$G_{T^{xy}T^{xy}}^{R}\left(t\right)\equiv\frac{4\delta^{2}W}{\delta h_{\mathrm{a}xy}\left(t\right)\delta h_{\mathrm{r}xy}\left(0\right)},$ (A.20)
where $W$ is the generating functional, $\delta h_{xy}$ denotes the shear perturbation of the metric, and the subscripts $\left(\mathrm{r,a}\right)$ indicate the closed time-path formalism.
Using the holographic dictionary, one can read the retarded correlator from eq. (A.12):
$G_{T^{xy}T^{xy}}^{R}\left(\omega\right)=L^{(3)}+3\frac{h^{(3)}}{h^{(0)}},$ (A.21)
@@ Appendix B Training scheme and report
There are various hyperparameter tuning methods, which consume different computing resources. Here we propose a simple way to quickly find the exponent $b$ in the square of graviton mass. We define the search range as $[0,4]$ with the step size 1. We use the standard cross-validation method with the simple hold-out technique Arlot . We separate the dataset within the frequency range $[0.01,\omega_{\max}]$ into two parts. The part with $[0.01,\omega_{\max}-1]$ goes to the training set and the part with $[\omega_{\max}-0.99,\omega_{\max}]$ to the validation set. The rationale for this separation lies in the dependence of the exponent $b$ on the UV physics. We use the optimizers RMSProp and Adam in order. Their learning rates, epoch numbers, and batch sizes are set as (0.001, 0.0001), (10, 10) and (1, 1). We also add the L1 regularization term $L_{1}=0.1\times\left|\theta\right|$ in the loss function. For each $b$ in the search range, we train 5 times and take the average validation error to select the optimal $b$. As long as $\omega_{\max}$ is large enough, the correct $b$ can be found for the current three targets. Specifically, $\omega_{\max}=2$ is large enough for the MG and LA models, and $\omega_{\max}=4$ is large enough for the GA model, see figure 6.
Figure 6: The hyperparameter tuning of $b$ in the MG (left), LA (middle) and GA (right) models. Each $b$ is trained five times, yielding five nearly coincident loss function points in most cases. As one can see, the optimal hyperparameters for three models are 1, 2, and 2, respectively.